# On dense orbits in the space of subequivalence relations

### v2

@@ 4.3 Dense orbits in the space of hyperfinite subequivalence relations
The equivalence of the first three items (i), (ii) and (iii) is a direct consequence of Theorem 4.7. The implication (i)$\implies$(iv) is exactly Theorem 4.11, and (iv) clearly implies (v). Finally, let us prove (v)$\implies$(iii): assume the closure of the $\mathrm{Aut}(\mathcal{R})$-orbit of $\mathcal{S}$ contains $\operatorname{\mathrm{Sub}}_{hyp}(\mathcal{R})$. Since $\Delta_{X}$ is hyperfinite, we have that $\Delta_{X}$ belongs to the closure of the $\mathrm{Aut}(\mathcal{R})$-orbit of $\mathcal{S}$, so we only have to show $\mathcal{S}$ is aperiodic. Assume by contradiction that $\mathcal{S}$ is not aperiodic, then for some $n\in\mathbb{N}$ we have $\mu(\{x\in X\colon\left\lvert[x]_{\mathcal{S}}\right\rvert\leqslant n\})>0$. Let $\delta=\mu(\{x\in X\colon\left\lvert[x]_{\mathcal{S}}\right\rvert\leqslant n\})$, then $\mathcal{S}$ belongs to the $\mathrm{Aut}(\mathcal{R})$-invariant set $\mathbf{B}$ of all $\mathcal{S}^{\prime}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ such that
@@
    $\mathcal{S}$ is aperiodic and the closure of the $[\mathcal{R}_{0}]$-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
    The $[\mathcal{R}_{0}]$-orbit of $\mathcal{S}$ is dense in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$;
@@
In the non-singular ergodic type II or type III case, one can show that there are always dense orbits in the space of subequivalence relations. Let us sketch the proof: first enumerate a dense subset of $\operatorname{\mathrm{Sub}}(\mathcal{R})$ as $(\mathcal{S}_{n})$ and note that there is a sequence $(\varphi_{n})$ of elements of $[[\mathcal{R}]]$ such that for all $n$, $\mathrm{dom}\;\varphi_{n}=X$ but $X=\bigsqcup_{n}\operatorname{\mathrm{rng}}\varphi_{n}$. Then the equivalence relation $\mathcal{S}=\bigsqcup_{n}\varphi_{n}\cdot\mathcal{S}_{n}$ has a dense full group orbit, as one can see by approximating each $\varphi_{n}$ by a full group element in the $d_{M}$ metric.
@@ 4.4 Meagerness of full group orbits
which is biinvariant, complete and separable, thus endowing $[\mathcal{R}]$ with a Polish group topology. We will use without mention the well-known fact that if $M(\mathcal{S}\cap T)>1-\epsilon$, then there is $S\in[\mathcal{S}]$ such that $d_{u}(S,T)<\epsilon$ (to see this, note that $\mathcal{S}\cap T\in[[\mathcal{S}]]$, hence it can be extended to an element $S$ of the full group of $\mathcal{S}$ by Lemma 4.2).
@@
We leave it to the reader to check that $\prec$ is $[\mathcal{R}]$-invariant, meaning that if $\mathcal{S}\prec\mathcal{T}$ and $U_{1},U_{2}\in[\mathcal{R}]$, then $U_{1}\cdot\mathcal{S}\prec U_{2}\cdot\mathcal{T}$. Also note that we always have $\mathcal{S}\prec\mathcal{S}$.
@@
Let $\mathcal{R}$ be an ergodic p.m.p. equivalence relation, consider the $[\mathcal{R}]$-action on the Polish space $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ of hyperfinite subequivalence relations of $\mathcal{R}$. Then all $[\mathcal{R}]$-orbits in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ are meager.
@@
Assume by contradiction there is a non meager $[\mathcal{R}]$-orbit. Since there is a dense $[\mathcal{R}]$-orbit in $\operatorname{\mathrm{Sub}}(\mathcal{R})$, the topological 00-$1$ law yields that there is a comeager orbit. Denoting by $E$ the equivalence relation generated by the $[\mathcal{R}]$-action on $\operatorname{\mathrm{Sub}}(\mathcal{R})$, we deduce that $E$ is comeager in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})\times\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$.
@@
Since $\prec$ is $[\mathcal{R}]$-invariant, its complement also is. This complement is moreover $G_{\delta}$ by Lemma 4.15, and by the above claim it contains the subset $[\mathcal{R}]\cdot\mathcal{S}\times[\mathcal{R}]\cdot\mathcal{T}$, which is dense in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})\times\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ by Theorem 4.11. So the complement of $\prec$ is comeager, and it should thus intersect $E$. This means that one can find a subequivalence relation $\mathcal{S}$ and $T\in[\mathcal{R}]$ such that $\mathcal{S}\not\prec T\cdot\mathcal{S}$, a contradiction. ∎
@@
Let $\mathcal{R}_{0}$ be the hyperfinite ergodic p.m.p. equivalence relation. Then all the $[\mathcal{R}_{0}]$-orbits are meager in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$
@@ 5 The uniform metric and complexity calculations
Suppose $\mathcal{R}$ is an aperiodic non-singular equivalence relation. There is a continuous reduction from the set of finite subsets of $\mathbb{N}$ to the complement of the set of subequivalence relations which are of infinite index, and to the set of finite index subequivalence relations, both for the uniform metric.
@@
Denote by $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})$ the space of finite index subequivalence relations. Note that by ergodicity, if $E$ has only finitely many ergodic components, then the space of finite index subequivalence relations is equal to the reunion over $k\in\mathbb{N}$ of the spaces $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ of all subequivalence relations whose index is uniformly less than $k$.
Let us show that for each $k\in\mathbb{N}$, the set $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is closed. Suppose $\mathcal{S}\notin\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$, then there is a positive measure set of $x\in X$ such that $[x]_{\mathcal{R}}$ contains at least $k+1$ distinct $\mathcal{S}$ classes. We thus find $A\subseteq X$ of positive measure and $\varphi_{1},...\varphi_{k}\in[[\mathcal{R}]]$ with common domain $A$ and such that $A,\operatorname{\mathrm{rng}}\varphi_{1},\dots,\operatorname{\mathrm{rng}}\varphi_{k}$ are and $\mathcal{S}$ is disjoint from the reunion of the graphs of the $\varphi_{i}\varphi_{j}^{-1}$. In other words we find a finite subequivalence relation of $\mathcal{R}$ which is disjoint from $\mathcal{S}$ (after of course removing $\Delta_{X}$) and all whose non trivial classes have cardinal $\geqslant k+1$.
@@
[…]
for $n$ large enough there will be a positive measure set of $x\in A$ such that for all $i\in\{1,\dots,k\}$, $(x,\varphi_{i}(x))\notin\mathcal{S}_{n}$, in particular $\mathcal{S}_{n}$ has somewhere index $>k$. So $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is closed and hence $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})=\bigcup_{k\in\mathbb{N}}\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is $F_{\sigma}$. Moreover, $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})=\bigcup_{k\in\mathbb{N}}\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is $F_{\sigma}$-complete by virtue of the preceding proposition.
Now if $\mathcal{R}$ has infinitely many ergodic components, let $(X_{n})$ be a partition of $X$ into $\mathcal{R}$-invariant sets of positive measure. Denote by $\mathcal{P}_{f}(\mathbb{N})\subseteq\{0,1\}^{\mathbb{N}}$ the set of finite subsets of $\mathbb{N}$, viewed as finitely supported functions $\mathbb{N}\to\{0,1\}$.
For every $n\in\mathbb{N}$, let $\Phi_{n}:\mathcal{\to}\operatorname{\mathrm{Sub}}(\mathcal{R}_{\restriction X_{n}})$ be the reduction from the previous proposition, and for a sequence $(x_{n})$ of elements of $\{0,1\}^{\mathbb{N}}$, let $\Phi(x_{n})=\bigsqcup_{n}\Phi_{n}(x_{n})\in\operatorname{\mathrm{Sub}}(\mathcal{R})$. By Lemma 3.11 $\Phi$ is continuous, and by construction $\Phi^{-1}(\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R}))=\prod_{n\in\mathbb{N}}\mathcal{P}_{f}(\mathbb{N})$. Being a (countable) infinite product of $F_{\sigma}$-hard sets, the latter is $F_{\sigma\delta}$-hard, so the set of finite index subequivalence relations is $F_{\sigma\delta}$-hard. ∎
@@
The following lemma is probably well-known and true as well in the non-singular case, but for simplicity we stick to the p.m.p. setup.
Lemma 5.6 (Dye).
For every aperiodic p.m.p. equivalence relation $\mathcal{R}$, there is a free $\Gamma_{0}$-action which induces a subequivalence relation of $\mathcal{R}$.
@@
By Maharam’s lemma, we may build by induction a partition $(A_{n})_{n\geqslant 0}$ of $X$ into subsets such that for all $n\in\mathbb{N}$, their conditional measure satisfies $\mu_{\mathcal{R}}(A_{n})=\frac{1}{2^{n+1}}$. For $n\in\mathbb{N}$, let now $B_{n}=X\setminus\bigsqcup_{m\leqslant n}A_{m}$, which also satisfies $\mu_{\mathcal{R}}(A_{n})=\frac{1}{2^{n+1}}$. Moreover, for every $n$ we have $B_{n}=A_{n+1}\sqcup B_{n+1}$ by construction.
By Lemma 4.2, for every $n\in\mathbb{N}$ there is $\varphi_{n}\in[[\mathcal{R}]]$ such that $\dim\varphi_{n}=A_{n}$ and $\operatorname{\mathrm{rng}}\varphi_{n}=B_{n}$. We will extend these to involutions so as to have a free $\Gamma_{0}$-action. Letting $G_{n}=\{f\in\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}\colon\forall m>n,f(m)=0\}$, we can write $\Gamma_{0}$ as the increasing union $\Gamma_{0}=\bigcup_{n}G_{n}$, and we have $G_{n}=\left\langle s_{0},\dots,s_{n}\right\rangle$, where $s_{i}(k)=0$ if $i\neq k$, and $s_{i}(k)=1$ if $i=k$. Our inductive construction of the action will be so that for all $n$,
$X=\bigsqcup_{g\in G_{n}}g\cdot B_{n}$ (3)
We begin by letting $s_{0}\cdot x=\varphi_{0}(x)$ if $x\in A_{0}$, and $s_{0}\cdot x=\varphi_{0}^{-1}(x)$ if $x\in B_{0}=X\setminus A_{0}$, then $X=B_{0}\sqcup s_{0}\cdot B_{0}$. And then, assuming that the action of $G_{n}$ has been defined so that $X=\bigsqcup_{g\in G_{n}}g\cdot B_{n}$, we begin by defining the $s_{n+1}$ action on $B_{n}$ by letting $s_{n+1}\cdot x=\varphi_{n+1}(x)$ if $x\in A_{n+1}$ and $s_{n+1}(x)=\varphi_{n+1}^{-1}(x)$ if $x\in B_{n+1}$. Since we want $s_{n+1}$ to commute with $G_{n}$ and we have $X=\bigsqcup_{g\in G_{n}}g\cdot B_{n}$, we are then forced to extend $s_{n+1}$ to the whole of $x$ by letting, for every $x\in B_{n}$ and every $g\in G_{n}$,
@@
It is then not hard to check that this does define a $G_{n+1}$ action with $X=\bigsqcup_{g\in G_{n+1}}g\cdot B_{n+1}$, which provides us a $\Gamma_{0}$-action by induction in the end. This action is free since Equation 3 implies that the action restricted to each $G_{n}$ is free, which finishes the proof. ∎
@@
It follows from [Dye59, Thm. 3] that when $\mathcal{R}$ is p.m.p. hyperfinite and aperiodic, one can even ask that $\mathcal{R}$ is equal to the equivalence relation generated by the $\Gamma_{0}$-action. This led Dye to formulate the question wether any p.m.p. equivalence relation should come the free action of some countable group (see the paragraph right before Lemma 6.5 in [Dye59]). This question was answered in the negative 40 years later by Furman [furmanOrbitEquivalenceRigidity1999].
@@
Let $\mathcal{R}$ be an aperiodic p.m.p. equivalence relation. Then $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathcal{R})$ is $F_{\sigma\delta}$-complete.
@@
Partition $X$ into $(X_{n})$ where each $X_{n}$ has positive measure. The restriction of $\mathcal{R}$ to each $X_{n}$ is aperiodic, by the previous lemma it thus contains a subequivalence relation coming from a free $\Gamma_{0}$-action $\alpha_{n}$. Now consider the map $\prod_{n}\mathrm{Sub}(\Gamma_{0})\to S(E)$ which takes any sequence $(\lambda_{n})$ to the equivalence relation $\bigsqcup_{n}\mathcal{R}_{\alpha_{n}(\Lambda_{n})}$. This is a topological embedding by Lemma 3.11, and it reduces the $F_{\sigma\delta}$-complete set $\operatorname{\mathrm{Sub}}_{\mathrm{fin}}(\Gamma_{0})^{\mathbb{N}}$ to the set of finite equivalence relations, which is thus $F_{\sigma\delta}$-hard. Since by [Kec, Thm. 8.5], $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathcal{R})$ is $F_{\sigma\delta}$, it is $F_{\sigma\delta}$-complete as wanted. ∎

### v3

@@ 4.3 Dense orbits in the space of hyperfinite subequivalence relations
[…]
@@
    $\mathcal{S}$ is aperiodic and the closure of the $\mathcal{R}_{0}]-$orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
    The $\mathcal{R}_{0}]-$orbit of $\mathcal{S}$ is dense in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$;
@@
In the non-singular ergodic type II or type III case, one can show that there are always dense orbits in the space of subequivalence relations. Let us sketch the proof: first note that by asumption there is a sequence $(\varphi_{n})$ of elements of $\mathcal{R}$]] such that for all $n$, $\mathrm{dom}\;\varphi_{n}=X$ but $X=\bigsqcup_{n}\operatorname{\mathrm{rng}}\varphi_{n}$ (this actually characterizes equivalence relations of type II or III among ergodic non-singular equivalence relations). We then enumerate a dense subset of $\operatorname{\mathrm{Sub}}(\mathcal{R})$ as $(\mathcal{S}_{n})$. The desired subequivalence relation with a dense orbit is $\mathcal{S}=\bigsqcup_{n}\varphi_{n}\cdot\mathcal{S}_{n}$, as one can see by approximating each $\varphi_{n}$ by a full group element in the $d_{M}$ metric.
@@ 4.4 Meagerness of full group orbits
which is biinvariant, complete and separable, thus endowing $\mathcal{R}$] with a Polish group topology. We will use without mention the well-known fact that if $M(\mathcal{S}\cap T)>1-\epsilon$, then there is $S\in[\mathcal{S}]$ such that $d_{u}(S,T)<\epsilon$ (to see this, note that $\mathcal{S}\cap T\in[[\mathcal{S}]]$, hence it can be extended to an element $S$ of the full group of $\mathcal{S}$ by Lemma 4.2).
@@
We will crucially use Ioana’s version of Popa’s interwining theorem [Ioa12, Lem. 1.7]: if $\mathcal{S}\prec\mathcal{T}$ then $\mathcal{S}$ can be translated by an element of the full group of $\mathcal{R}$ so as to have somewhere finite index in $\mathcal{T}$. The interested reader is also refered to [Spa23, Lem. 3.1] for a characterization of $\mathcal{S}\prec\mathcal{T}$ when $\mathcal{S}$ has infinite index in $\mathcal{R}$.
We leave it to the reader to check that $\prec$ is $\mathcal{R}$]-invariant, meaning that if $\mathcal{S}\prec\mathcal{T}$ and $U_{1},U_{2}\in[\mathcal{R}]$, then $U_{1}\cdot\mathcal{S}\prec U_{2}\cdot\mathcal{T}$. Also note that we always have $\mathcal{S}\prec\mathcal{S}$.
@@
Let $\mathcal{R}$ be an ergodic p.m.p. equivalence relation, consider the $\mathcal{R}$]-action on the Polish space $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ of hyperfinite subequivalence relations of $\mathcal{R}$. Then all $\mathcal{R}$]-orbits in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ are meager.
@@
Assume by contradiction there is a non meager $\mathcal{R}$]-orbit. Since there is a dense $\mathcal{R}$]-orbit in $\operatorname{\mathrm{Sub}}(\mathcal{R})$, the topological 0-11 law yields that there is a comeager orbit. Denoting by $E$ the equivalence relation generated by the $\mathcal{R}$]-action on $\operatorname{\mathrm{Sub}}(\mathcal{R})$, we deduce that $E$ is comeager in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})\times\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$.
@@
Since $\prec$ is $\mathcal{R}$]-invariant, its complement also is. This complement is moreover $G_{\delta}$ by Lemma 4.15, and by the above claim it contains the subset $\mathcal{R}]\cdot\mathcal{S}\times[\mathcal{R}]\cdot\mathcal{T}$, which is dense in $\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})\times\operatorname{\mathrm{Sub}_{\mathrm{hyp}}}(\mathcal{R})$ by Theorem 4.11. So the complement of $\prec$ is comeager, and it should thus intersect $E$. This means that one can find a subequivalence relation $\mathcal{S}$ and $T\in[\mathcal{R}]$ such that $\mathcal{S}\not\prec T\cdot\mathcal{S}$, a contradiction. ∎
@@
Let $\mathcal{R}_{0}$ be the hyperfinite ergodic p.m.p. equivalence relation. Then all the $\mathcal{R}_{0}]-$orbits are meager in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$
@@ 5 The uniform metric and complexity calculations
In what follows, we freely identify the Cantor space $\{0,1\}^{\mathbb{N}}$ to the set of subsets of $\mathbb{N}$, and the continuous reduction alluded to is thus defined on the whole $\{0,1\}^{\mathbb{N}}$.
@@
Suppose $\mathcal{R}$ is an aperiodic non-singular equivalence relation. There is a continuous reduction from the set $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N})$ of finite subsets of $\mathbb{N}$ to the complement of the set of subequivalence relations which are of infinite index, and to the set of finite index subequivalence relations, both for the uniform metric.
@@
Denote by $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})$ the space of finite index subequivalence relations. Note that by ergodicity, if $E$ has only finitely many ergodic components, then the space of finite index subequivalence relations is equal to the union over $k\in\mathbb{N}$ of the spaces $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ of all subequivalence relations whose index is uniformly less than $k$.
Let us show that for each $k\in\mathbb{N}$, the set $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is closed. Suppose $\mathcal{S}\notin\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$, then there is a positive measure set of $x\in X$ such that $x]_{\mathcal{R}}$ contains at least $k+1$ distinct $\mathcal{S}$ classes. We thus find $A\subseteq X$ of positive measure and $\varphi_{1},...\varphi_{k}\in[[\mathcal{R}]]$ with common domain $A$ and such that $A,\operatorname{\mathrm{rng}}\varphi_{1},\dots,\operatorname{\mathrm{rng}}\varphi_{k}$ are and $\mathcal{S}$ is disjoint from the union of the graphs of the $\varphi_{i}\varphi_{j}^{-1}$. In other words we find a finite subequivalence relation of $\mathcal{R}$ which is disjoint from $\mathcal{S}$ (after of course removing $\Delta_{X}$) and all whose non trivial classes have cardinal $\geqslant k+1$.
@@
$M(\varphi_{1}\sqcup\cdots\sqcup\varphi_{k}\cap\mathcal{S}_{n})=\int_{A}\left\lvert\{i\in\{1,\dots,k\}\colon(x,\varphi_{i}(x))\in\mathcal{S}_{n}\}\right\rvert d\mu(x)\to 0,$
for $n$ large enough there will be a positive measure set of $x\in A$ such that for all $i\in\{1,\dots,k\}$, $(x,\varphi_{i}(x))\notin\mathcal{S}_{n}$, in particular $\mathcal{S}_{n}$ has somewhere index >k>k. So $\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is closed and hence $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})=\bigcup_{k\in\mathbb{N}}\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is $F_{\sigma}$. Moreover, $\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R})=\bigcup_{k\in\mathbb{N}}\operatorname{\mathrm{Sub}}_{[\leqslant k]}(\mathcal{R})$ is $F_{\sigma}$-complete by virtue of the preceding proposition.
Now if $\mathcal{R}$ has infinitely many ergodic components, let $(X_{n})$ be a partition of $X$ into $\mathcal{R}$-invariant sets of positive measure. Denote by $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N})\subseteq\{0,1\}^{\mathbb{N}}$ the set of finite subsets of $\mathbb{N}$, viewed as finitely supported functions $\mathbb{N}\to\{0,1\}$.
For every $n\in\mathbb{N}$, let $\Phi_{n}:\{0,1\}^{\mathbb{N}}\to\operatorname{\mathrm{Sub}}(\mathcal{R}_{\restriction X_{n}})$ be the continuous reduction from Proposition 5.2, and for a sequence $(x_{n})_{n\in\mathbb{N}}$ of elements of $\{0,1\}^{\mathbb{N}}$, let
$\Phi((x_{n})_{n\in\mathbb{N}})=\bigsqcup_{n\in\mathbb{N}}\Phi_{n}(x_{n})\in\operatorname{\mathrm{Sub}}(\mathcal{R}).$
By Lemma 3.11 $\Phi$ is continuous, and by construction
$\Phi^{-1}(\operatorname{\mathrm{Sub}}_{[<\infty]}(\mathcal{R}))=\prod_{n\in\mathbb{N}}\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N}).$
Being a (countable) infinite product of $F_{\sigma}$-hard sets, the latter is $F_{\sigma\delta}$-hard, so the set of finite index subequivalence relations is $F_{\sigma\delta}$-hard. ∎
@@
The following proposition was stated and proven in the p.m.p. setup in the first version of this paper (in this restricted setup, it is a consequence of [Dye59, Thm. 4]). We are very grateful to the referee for pointing out that the same result is true in the purely Borel setup and for indicating the proof.
Proposition 5.6.
For every aperiodic CBER $\mathcal{R}$, there is a free $\Gamma_{0}$-action which induces a subequivalence relation of $\mathcal{R}$.
We rely on the following key lemma.
Lemma 5.7.
Let $\mathcal{R}$ be an aperiodic CBER, then its Borel full group contains a fixed-point free involution.
@@
By [KM04, Lem. 7.4], $\mathcal{R}$ contains a Borel subequivalence relation $\mathcal{S}$, all whose classes have cardinality $2$. Our involution $U$ is then defined by: for all $x\in X$, $U(x)$ is the only $y\neq x$ such that $(x,y)\in\mathcal{S}$. ∎
Proof of Proposition 5.6.
We first inductively define a decreasing sequence $(A_{n})_{n\in\mathbb{N}}$ of Borel subsets of $X$ and fixed point-free partial involutions $(\varphi_{n})_{n\in\mathbb{N}}$ in the Borel pseudo full group of $\mathcal{R}$ with support $A_{n}$ as follows. We begin with an everywhere defined involution $\varphi_{0}$ with support $A_{0}=X$ as provided by the previous lemma and then, assuming $\varphi_{n}$ has been built, we let $A_{n+1}$ be a Borel fundamental domain for $\varphi_{n}$. Noting that the restriction of $\mathcal{R}$ to $A_{n+1}$ has to remain aperiodic, we use the above lemma to find $\varphi_{n+1}$ as a fixed point-free involution in the Borel full group of the restriction of $\mathcal{R}$ to $A_{n+1}$.
We will extend these to involutions so as to have a free $\Gamma_{0}$-action. Letting $G_{n}=\{f\in\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}\colon\forall m>n,f(m)=0\}$, we can write $\Gamma_{0}$ as the increasing union $\Gamma_{0}=\bigcup_{n}G_{n}$, and we have $G_{n}=\left\langle s_{0},\dots,s_{n}\right\rangle$, where $s_{i}(k)=0$ if $i\neq k$, and $s_{i}(k)=1$ if $i=k$. Our inductive construction of the action will be so that for all $n$,
$X=\bigsqcup_{g\in G_{n}}g\cdot A_{n+1}$ (3)
We begin by letting $s_{0}\cdot x=\varphi_{0}(x)$ for all $x\in X$. And then, assuming that the action of $G_{n}$ has been defined so that $X=\bigsqcup_{g\in G_{n}}g\cdot A_{n+1}$, we start defining the $s_{n+1}$ action on $A_{n+1}$ by letting $s_{n+1}\cdot x=\varphi_{n+1}(x)$ if $x\in A_{n+1}$. Since we want $s_{n+1}$ to commute with $G_{n}$ and we have $X=\bigsqcup_{g\in G_{n}}g\cdot A_{n+1}$, we are then forced to extend $s_{n+1}$ to the whole of $x$ by letting, for every $x\in A_{n+1}$ and every $g\in G_{n}$,
@@
It is then not hard to check that this does define a $G_{n+1}$ action with $X=\bigsqcup_{g\in G_{n+1}}g\cdot A_{n+2}$, which provides us a $\Gamma_{0}$-action by induction in the end. This action is free since Equation (3) implies that the action restricted to each $G_{n}$ is free, which finishes the proof. ∎
@@
It follows from [Dye59, Thm. 3] that when $\mathcal{R}$ is p.m.p. hyperfinite and aperiodic, one can upgrade the above result and obtain that $\mathcal{R}$ is equal to the equivalence relation generated by the $\Gamma_{0}$-action. This led Dye to formulate the question whether any p.m.p. equivalence relation should come the free action of some countable group (see the paragraph right before Lemma 6.5 in [Dye59]). This question was answered in the negative 40 years later by Furman [Fur99].
@@
Let $\mathcal{R}$ be an aperiodic non-singular equivalence relation. Then $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathcal{R})$ is $F_{\sigma\delta}$-complete.
@@
Partition $X$ into $(X_{n})_{n\in\mathbb{N}}$ so that each $X_{n}$ has positive measure and the restriction of $\mathcal{R}$ to each $X_{n}$ is aperiodic. By the previous lemma each restriction of $\mathcal{R}$ to $X_{n}$ contains a subequivalence relation coming from a free $\Gamma_{0}$-action $\alpha_{n}$. Now consider the map $(\mathrm{Sub}(\Gamma_{0}))^{\mathbb{N}}\to\operatorname{\mathrm{Sub}}(\mathcal{R})$ which takes any sequence $(\lambda_{n})_{n\in\mathbb{N}}$ to the equivalence relation $\bigsqcup_{n\in\mathbb{N}}\mathcal{R}_{\alpha_{n}(\Lambda_{n})}$. This is a topological embedding by Lemma 3.11, and it reduces the $F_{\sigma\delta}$-complete set $\operatorname{\mathrm{Sub}}_{\mathrm{fin}}(\Gamma_{0})^{\mathbb{N}}\subseteq\operatorname{\mathrm{Sub}}(\Gamma_{0})^{\mathbb{N}}$ to the set of finite equivalence relations, which is thus $F_{\sigma\delta}$-hard.
We are left with proving that $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathcal{R})$ is $F_{\sigma\delta}$, which is due to Kechris in the p.m.p. setup [Kec, Thm. 8.5]. While his proof generalizes to the non-singular setup, we provide here another approach relying on spaces of measurable maps. Let us fix a uniquely generating sequence of moving partial involutions $(\varphi_{n})_{n\in\mathbb{N}}$, consider for every $n\in\mathbb{N}$ the continuous map
$\iota_{n}:\operatorname{\mathrm{Sub}}(\mathcal{R})\to\mathrm{MAlg}(X,\mu)$
which associates to every $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ the set of $x\in\mathrm{dom}\;\varphi_{n}$ such that $(x,\varphi_{n}(x))\in\mathcal{S}$.
Identifying subsets to characteristic functions yields a natural homeomorphism $\Phi$ between $\mathrm{MAlg}(X,\mu)$ and the Polish space of measurable maps $\mathrm{L}^{0}(X,\mu,\{0,1\})$ (see [LMS25, Appendix B] for more on $\mathrm{L}^{0}$ spaces). Using this homeomorphism and putting all the maps $\iota_{n}$ together, we get an continuous embedding
$\iota=(\Phi\circ\iota_{n})_{n\in\mathbb{N}}:\operatorname{\mathrm{Sub}}(\mathcal{R})\hookrightarrow\prod_{n\in\mathbb{N}}\mathrm{L}^{0}(X,\mu,\{0,1\})\simeq\mathrm{L}^{0}(X,\mu,\{0,1\}^{\mathbb{N}}).$
We will conclude the proof via a straightforward application of the claim below.
Claim.
Let $Y$ be a Polish space, let $D\subseteq Y$ be $F_{\sigma}$. Then $\mathrm{L}^{0}(X,\mu,D)$ is an $F_{\sigma\delta}$ subset of $\mathrm{L}^{0}(X,\mu,Y)$.
Proof of the claim.
It follows directly from the definition of the topology of convergence in measure given in [LMS25, Appendix B] that for any open subset $U\subseteq Y$, the map $f\in\mathrm{L}^{0}(X,\mu,Y)\mapsto\mu(f^{-1}(U))$ is lower semi-continuous. Taking complements, given a closed subset $C\subseteq\{0,1\}^{\mathbb{N}}$, the map $f\in\mathrm{L}^{0}(X,\mu,Y)\mapsto\mu(f^{-1}(C))$ is thus upper semi-continuous. Now write $D=\bigcup_{n}C_{n}$ as an increasing union of closed subsets $C_{n}$, and observe that $\mu(f^{-1}(D)))=1$ iff for all $\epsilon>0$ there is $n$ such that $\mu(f^{-1}(C_{n}))\geqslant 1-\epsilon$. ∎
In order to finish the proof, take $D=\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N})$ to be the $F_{\sigma}$ set of finite subsets of $\mathbb{N}$, viewed as a subset of $Y=\{0,1\}^{\mathbb{N}}$. Then $\mathrm{L}^{0}(X,\mu,\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N}))$ is $F_{\sigma\delta}$ in $\mathrm{L}^{0}(X,\mu,\{0,1\}^{\mathbb{N}})$ by the above claim. Finally, observe that $\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathcal{R})=\iota^{-1}(\mathrm{L}^{0}(X,\mu,\operatorname{\mathrm{Sub}_{\mathrm{fin}}}(\mathbb{N})))$ which is thus $F_{\sigma\delta}$ by continuity of $\iota$. ∎